# PolaRiS: Scalable Real-to-Sim Evaluations for Generalist Robot Policies

### v1

@@ 5.1 Experimental Setup
Robot platform. We perform our experiments using the DROID [21] platform–one of the mostly widely used robot platforms for generalist policies [56, 2, 22]–which consists of a 7-DoF Franka Panda robot arm, and a set of two ZED RGB cameras with one mounted to the robot’s wrist. While DROID offers both joint velocity and joint position action spaces, we focus on evaluating policies using joint position actions, which we found easier to simulate accurately since they suffer from less accumulating error.
Policies. We use a representative set of DROID VLAs from prior works for testing the efficacy of PolaRiS as an evaluation tool: (1) $\bm{\pi_{0}}$ [1] and $\bm{\pi_{0}}$-FAST [56], which are diffusion-based and auto-regressive VLAs respectively, (2) PaliGemma-binning, an auto-regressive VLA with binning action tokenization from [56], and (3) $\bm{\pi_{0.5}}$ [2], the current state-of-the-art VLA on DROID [22].
@@
Comparisons. We compare our real-to-sim evaluation approach to existing methods used for offline policy evaluation: MSE, which computes training / validation set action MSE for each policy; Libero-Score, which evaluates policies on the Libero simulation benchmark [11] that is often used in VLA benchmarking [57, 58, 8, 59]; and Video model evaluation, which rolls out policies in the open-source Ctrl-World [29] video model and uses humans to score the rollouts. For Libero, to get the highest-quality evaluations possible, we finetune all policies on Libero-90, the largest of the Libero datasets, for 50k steps and evaluate on all 90 tasks, running 50 rollouts per task or a total of 4,500 evaluations per policy. Notably, this is a more thorough evaluation than done by most existing VLA papers using Libero, which typically evaluate on much smaller Libero splits [57, 58, 8]. Additionally, we evaluate multiple Libero-finetuned checkpoints (1k, 10k, 50k) and report results for the best-correlated checkpoint in Figure˜7. As a result, the Libero-Score approach uses approximately 10x more finetuning and evaluation compute than our simulated evaluation approach. We note that existing real-to-sim evaluation approaches [9, 34], including SIMPLER, do not support evaluation of policies with wrist cameras, and thus cannot even be used to evaluate most state of the art VLA models.
@@ 5.2 Real-to-Sim Correlation in Unseen Environments
Figure 8: Correlation with RoboArena [22]: PolaRiS evaluations show strong performance correlation with RoboArena [22] rankings.
We report real-world correlation results for all tested approaches in Figure˜7 and Figure˜7. In line with existing literature [9], we find that MSE is a poor metric for real-world policy performance and shows low correlation. Libero-score too is poorly correlated with real-world performance, with nearly all tested policies scoring between 90 and 95% success in Libero, but spanning the full spectrum from high-performing to low-performing policies in the real world. Finally, we find that today’s open-source video models are not yet reliable enough for policy evaluation in diverse scenes: in our experiments, Ctrl-World often produces heavy hallucinations during object interaction, making accurate scoring challenging and ultimately leading to clear policy mis-rankings (MMRV: 0.22). While recent works have shown promising video model evaluation results in narrower scene settings [29, 30], our experiments suggest that more advances are needed for these models to be widely applicable tools for robot policy evaluation in diverse real-world tasks.
@@
Furthermore, we find that PolaRiS scores exhibit strong correlation to RoboArena [22] real-world evaluation scores (see Figure˜8). We report correlation to the average progress scores reported as part of the RoboArena evaluations. While we find that scores correlate well, we emphasize that the simulated evaluations in our experiments only capture a small subset of the capabilities tested in RoboArena—particularly those involving non-rigid body dynamics, soft-body interactions, and complex contact behaviors that remain challenging for current simulators. Thus, we believe that PolaRiS can serve as a useful tool for cheaply and rapidly estimating VLA performance during research and development, but does not entirely replace real-world evaluation.
@@ 5.3 Ablation Studies and Analysis
Visual fidelity. To ablate the impact of the visual fidelity of simulated environments on real-to-sim correlation, we create versions of our simulated evaluation scenes with different levels of visual quality (see Table 5), using ray-tracing instead of Gaussian splatting, and removing background and foreground textures. We report real-to-sim correlation after co-training all policies with the respective simulated dataset in Figure˜11, left. First, we find that after co-training, all simulation variants can provide decently correlated results, suggesting that the simulation data co-training can alleviate some visual sensitivity observed in prior works [9]. Yet, we find that PolaRiS with Gaussian splatting rendering leads to the strongest correlation. While these results suggest visual fidelity is not the sole driver of high-quality simulated evaluations, the comparison to Libero-Score (Section˜5.2) underlines the benefit of real-to-sim evaluation techniques compared to hand-crafting low-fidelity simulated benchmarks: real-to-sim evaluation scans real-world environments and thus allows evaluation in environments that are closer to the training distribution.
@@ 6 Discussion
While a promising step for policy evaluation, there are still several points of improvement for PolaRiS. Going beyond Gaussian splat rendering to richer diffusion based renderings can further bridge the real-to-sim visual gaps. Moreover, PolaRiS performs a relatively simplistic identification of system dynamics. We can draw on a rich literature of system identification [60] to better identify the robot and environment dynamics. Lastly, we can consider hybrid simulation techniques, merging elements of world modeling with physics-based simulation to generate more scalable simulated evaluations, including more complex objects and scene configurations.
@@ Acknowledgments
This research was partially support by Amazon Science Hub, Army Research Lab, the DARPA TIAMAT Program, and the Toyota Research Institute University 2.0 Program, with compute support from Microsoft. The authors would like to thank Ludwig Schmidt, Masha Itkina, Junjie Ye, Jiawei Yang for the early discussion on the project, Mateo Guaman Castro for help and feedback on figures and presentation, and Apurva Badithela and Lihan Zha for feedback on an earlier draft of this paper. We also thank Aditya Shah and Daniel Gorbatov for helping with some evaluations.
@@ Appendix A Additional Details for Preliminaries
Here $\mathcal{R}_{\text{dist}}$ is a depth regularizer that encourages each disk to align closely along the viewing ray to reduce thickness and multi-view inconsistency, and $\mathcal{R}_{\text{norm}}$ is a normal regularizer that minimizes the discrepancy between disk normals and the gradient of the rendered depth map. Once optimized, the disks give rise to both high-quality novel-view synthesis and accurate geometry that can be extracted (e.g., via meshing) and used for simulation of robotic policies. While closely related to 3D Gaussian splatting [19], the 2DGS algorithm more faithfully captures object geometries while maintaining high fidelity visuals. As we show in our experimental analysis, this ensures better zero-shot evaluation performance for robot policies trained in the real world.
@@ B.1 Reconstruction
To ensure all reconstructions are expressed in a consistent world coordinate frame, we define the ChArUco board origin as the canonical base frame $\mathcal{F}_{0}$. For each video, we get recovered camera poses $\{T_{i}\}_{i=1}^{N}$ for almost all video images in an arbitrary frame from COLMAP, and a subset of these frames will have measured camera coordinates $\{C_{i}\}_{i=1}^{M};M\leq N$ with respect to the ChArUco board. Using the COLMAP’s model aligner functionality [51], we are able to transform the collective camera poses’ arbitrary scale and orientation into an accurately scaled and oriented set of camera poses in the ChArUco board canonical base frame.
@@ C.2 Cotraining
For all of our experiments involving cotraining policies, we collected data in simulation environments using a twin of the real DROID platform [21]. To restrict behaviors from deviating significantly from the original DROID dataset, we adopted the same data collection scheme that was used in DROID. Data in simulation was collected with an Oculus Quest 2, and we setup our environments such that they could interface directly with the VR controller code shipped with the DROID code release. After data is collected, it’s processed into the RLDS format, exactly matching the structure of the original DROID RLDS. Finally, policies were trained using the open-source DROID policy training code provided in openpi [61], with a few modifications in dataloading. All cotraining was done with a mixture of 90% real DROID data and 10% simulated DROID data. Additional hyperparameters for cotraining can be found in Table 2 and Table 3. For our analysis of how many finetuning steps is needed to achieve best correlations, we ran independent finetuning runs with unique seeds (1-5). For each checkpoint we do 50 independent rollouts with the same initial conditions. All cotraining (both simulation and real world) is done on joint position action data. We find joint position policies to transfer much better from real to sim, with the intuition of joint position policies avoiding compounding error. We demonstrate this intuition in Fig. 12 using open loop replay comparing joint velocity and joint position policies.

### v2

@@ 5.1 Experimental Setup
Robot platform. We perform our experiments using the DROID [khazatsky24droid] platform–one of the mostly widely used robot platforms for generalist policies [pertsch2025fast, pi05, atreya2025roboarena]–which consists of a 7-DoF Franka Panda robot arm, and a set of two ZED RGB cameras with one mounted to the robot’s wrist. While DROID offers both joint velocity and joint position action spaces, we focus on evaluating policies using joint position actions, which we found easier to simulate accurately since they suffer from less accumulating error.
Policies. We use a representative set of DROID VLAs from prior works for testing the efficacy of PolaRiS as an evaluation tool: (1) $\bm{\pi_{0}}$ [blackpi0] and $\bm{\pi_{0}}$-FAST [pertsch2025fast], which are diffusion-based and auto-regressive VLAs respectively, (2) PaliGemma-binning, an auto-regressive VLA with binning action tokenization from [pertsch2025fast], and (3) $\bm{\pi_{0.5}}$ [pi05], the current state-of-the-art VLA on DROID [atreya2025roboarena].
@@
Comparisons. We compare our real-to-sim evaluation approach to existing methods used for offline policy evaluation: MSE, which computes training / validation set action MSE for each policy; Libero-Score, which evaluates policies on the Libero simulation benchmark [liu23libero] that is often used in VLA benchmarking [huang25otter, kim25vla, kim24openvla, lee2025molmoactactionreasoningmodels]; and Video model evaluation, which rolls out policies in the open-source Ctrl-World [guo2025ctrl] video model and uses humans to score the rollouts. For Libero, to get the highest-quality evaluations possible, we finetune all policies on Libero-90, the largest of the Libero datasets, for 50k steps and evaluate on all 90 tasks, running 50 rollouts per task or a total of 4,500 evaluations per policy. Notably, this is a more thorough evaluation than done by most existing VLA papers using Libero, which typically evaluate on much smaller Libero splits [huang25otter, kim25vla, kim24openvla]. Additionally, we evaluate multiple Libero-finetuned checkpoints (1k, 10k, 50k) and report results for the best-correlated checkpoint in Figure˜7. As a result, the Libero-Score approach uses approximately 10x more finetuning and evaluation compute than our simulated evaluation approach. We note that existing real-to-sim evaluation approaches [li24simpler, jangir2025robotarenainftyscalablerobot], including SIMPLER, do not support evaluation of policies with wrist cameras, and thus cannot even be used to evaluate most state of the art VLA models.
@@ 5.2 Real-to-Sim Correlation in Unseen Environments
Figure 8: Correlation with RoboArena [atreya2025roboarena]: PolaRiS evaluations show strong performance correlation with RoboArena [atreya2025roboarena] rankings.
We report real-world correlation results for all tested approaches in Figure˜7 and Figure˜7. In line with existing literature [li24simpler], we find that MSE is a poor metric for real-world policy performance and shows low correlation. Libero-score too is poorly correlated with real-world performance, with nearly all tested policies scoring between 90 and 95% success in Libero, but spanning the full spectrum from high-performing to low-performing policies in the real world. Finally, we find that today’s open-source video models are not yet reliable enough for policy evaluation in diverse scenes: in our experiments, Ctrl-World often produces heavy hallucinations during object interaction, making accurate scoring challenging and ultimately leading to clear policy mis-rankings (MMRV: 0.22). While recent works have shown promising video model evaluation results in narrower scene settings [guo2025ctrl, veorobotics2025], our experiments suggest that more advances are needed for these models to be widely applicable tools for robot policy evaluation in diverse real-world tasks.
@@
Furthermore, we find that PolaRiS scores exhibit strong correlation to RoboArena [atreya2025roboarena] real-world evaluation scores (see Figure˜8). We report correlation to the average progress scores reported as part of the RoboArena evaluations. While we find that scores correlate well, we emphasize that the simulated evaluations in our experiments only capture a small subset of the capabilities tested in RoboArena—particularly those involving non-rigid body dynamics, soft-body interactions, and complex contact behaviors that remain challenging for current simulators. Thus, we believe that PolaRiS can serve as a useful tool for cheaply and rapidly estimating VLA performance during research and development, but does not entirely replace real-world evaluation.
@@ 5.3 Ablation Studies and Analysis
Visual fidelity. To ablate the impact of the visual fidelity of simulated environments on real-to-sim correlation, we create versions of our simulated evaluation scenes with different levels of visual quality (see Table 5), using ray-tracing instead of Gaussian splatting, and removing background and foreground textures. We report real-to-sim correlation after co-training all policies with the respective simulated dataset in Figure˜11, left. First, we find that after co-training, all simulation variants can provide decently correlated results, suggesting that the simulation data co-training can alleviate some visual sensitivity observed in prior works [li24simpler]. Yet, we find that PolaRiS with Gaussian splatting rendering leads to the strongest correlation. While these results suggest visual fidelity is not the sole driver of high-quality simulated evaluations, the comparison to Libero-Score (Section˜5.2) underlines the benefit of real-to-sim evaluation techniques compared to hand-crafting low-fidelity simulated benchmarks: real-to-sim evaluation scans real-world environments and thus allows evaluation in environments that are closer to the training distribution.
@@ 6 Discussion
While a promising step for policy evaluation, there are still several points of improvement for PolaRiS. Going beyond Gaussian splat rendering to richer diffusion based renderings can further bridge the real-to-sim visual gaps. Moreover, PolaRiS performs a relatively simplistic identification of system dynamics. We can draw on a rich literature of system identification [systemID] to better identify the robot and environment dynamics. Lastly, we can consider hybrid simulation techniques, merging elements of world modeling with physics-based simulation to generate more scalable simulated evaluations, including more complex objects and scene configurations.
@@ Acknowledgments
This research was partially support by Amazon Science Hub, Army Research Lab, the DARPA TIAMAT Program, and the Toyota Research Institute University 2.0 Program, with compute support from Microsoft. The authors would like to thank Alex Fang, Ludwig Schmidt, Masha Itkina, Junjie Ye, Jiawei Yang for the early discussion on the project, Mateo Guaman Castro for help and feedback on figures and presentation, and Apurva Badithela and Lihan Zha for feedback on an earlier draft of this paper. We also thank Aditya Shah and Daniel Gorbatov for helping with evaluations, and Jenai (Xuning) Yang for advice and support with simulation.
[…]
@@ Appendix A Additional Details for Preliminaries
Here $\mathcal{R}_{\text{dist}}$ is a depth regularizer that encourages each disk to align closely along the viewing ray to reduce thickness and multi-view inconsistency, and $\mathcal{R}_{\text{norm}}$ is a normal regularizer that minimizes the discrepancy between disk normals and the gradient of the rendered depth map. Once optimized, the disks give rise to both high-quality novel-view synthesis and accurate geometry that can be extracted (e.g., via meshing) and used for simulation of robotic policies. While closely related to 3D Gaussian splatting [3dgs], the 2DGS algorithm more faithfully captures object geometries while maintaining high fidelity visuals. As we show in our experimental analysis, this ensures better zero-shot evaluation performance for robot policies trained in the real world.
@@ B.1 Reconstruction
To ensure all reconstructions are expressed in a consistent world coordinate frame, we define the ChArUco board origin as the canonical base frame $\mathcal{F}_{0}$. For each video, we get recovered camera poses $\{T_{i}\}_{i=1}^{N}$ for almost all video images in an arbitrary frame from COLMAP, and a subset of these frames will have measured camera coordinates $\{C_{i}\}_{i=1}^{M};M\leq N$ with respect to the ChArUco board. Using the COLMAP’s model aligner functionality [schonberger2016colmap], we are able to transform the collective camera poses’ arbitrary scale and orientation into an accurately scaled and oriented set of camera poses in the ChArUco board canonical base frame.
@@ C.2 Cotraining
For all of our experiments involving cotraining policies, we collected data in simulation environments using a twin of the real DROID platform [khazatsky24droid]. To restrict behaviors from deviating significantly from the original DROID dataset, we adopted the same data collection scheme that was used in DROID. Data in simulation was collected with an Oculus Quest 2, and we setup our environments such that they could interface directly with the VR controller code shipped with the DROID code release. After data is collected, it’s processed into the RLDS format, exactly matching the structure of the original DROID RLDS. Finally, policies were trained using the open-source DROID policy training code provided in openpi [openpi], with a few modifications in dataloading. All cotraining was done with a mixture of 90% real DROID data and 10% simulated DROID data. Additional hyperparameters for cotraining can be found in Table 2 and Table 3. For our analysis of how many finetuning steps is needed to achieve best correlations, we ran independent finetuning runs with unique seeds (1-5). For each checkpoint we do 50 independent rollouts with the same initial conditions. All cotraining (both simulation and real world) is done on joint position action data. We find joint position policies to transfer much better from real to sim, with the intuition of joint position policies avoiding compounding error. We demonstrate this intuition in Fig. 12 using open loop replay comparing joint velocity and joint position policies.